40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Entry  Tue Dec 14 14:10:36 2021, aaron, Summary, PSOMA, PSOMA calibration parameters, minor modifications 
    Reply  Fri Dec 17 12:46:41 2021, aaron, Summary, PSOMA, PSOMA calibration parameters, minor modifications 
       Reply  Fri Dec 17 18:46:55 2021, aaron, Summary, PSOMA, PSOMA calibration parameters, minor modifications 
Message ID: 2834     Entry time: Tue Dec 14 14:10:36 2021     Reply to this: 2836
Author: aaron 
Type: Summary 
Category: PSOMA 
Subject: PSOMA calibration parameters, minor modifications 

labwork to do

  • posting table of important parameters (like those in elog 2830) in wiki, to be printed and posted in the lab
  • psoma servo drawing

done

  • AC power for (most) electronics involved in locking share a single power strip. The laser drivers, LB box, function generator (connected to LB box), photodiodes supply, oscilloscope, spectrum analyzer (HP 8560E), and moku all are now powered from the long power strip ('A') located under the optics table. The OCXO preamplifier box supplying ~30 MHz modulations for PDH is powered by the Sorensen DC supplies... which are not powered from the same strip. The Sorensens are on the Aux rack, while the other electronics are on the PSOMA rack. It would be somewhat involved to power the Sorensens from the same supply (need a longer power cable sent across the walkway), but can be done.
  • I tied the optical table to ground through a long (springy) wire with 1 MOhm resistor soldered inline near the halfway point. One end of the wire is bolted to the optical table, and the other is bolted to the enclosure frame. There are two power strips bolted to the enclosure... but the 'enclosure' power strips are plugged in to the wall on the opposite side of the room from the 'under the table' power strip powering the PDH electronics. That's not great. I could move the PDH electronics over to the 'enclosure' power strips, but it would require some cable rerouting.

notes

  • The DB9 cable connecting the North TEC driver to the laser is flaky (moving the cable causes the TEC alarm to trip on the driver). The rubber of the cable is visibly separated from the casing, and the ground shielding is exposed no. I have powered down the North laser, and suggest we don't run it until replacing or repairing the cable
ELOG V3.1.3-